OpenStudy (anonymous):

Solve the following system of linear equations algebraically -2x+y=2 x+y=-1

OpenStudy (tux):

x=-1-y In equation -2x+y=2 substitute x=-1-y -2*(-1-y)+y=2 2+2y+y=2+2+3y=2 3y=0 y=0 go to x=-1-y and substitute y=0. Then you get x=-1